• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 08:04:13 on April28,2007

    Entry number 200369

    keyword=Owl Shift Summary

    00:00 Start shift with run 1916 in progress.

    02:24 Beam is gone, Stopped run 1921.

    02:32 Called MCC: one of the cavities had a problem. Problem fixed and beam is coming back.

    04:32 Move target to Dummy.

    04:41 Beam back. BPMAY position is consistently at 1.4 mm. Talked with MCC to fix it. BPMAY position was being tuned during run 2926.

    05:22 BPMAY position is now at 1.0 mm as requested.

    07:26 Move target to07:26 Move target to 4He.

    07:35 Asked MCC for 100 uA on 4He target to continue production runs.

    07:43 BPMA and BPMB y positions finally stabilize at 1 mm.
    4He.



    Run Summary
    -------------

    Run # Begin End Comments
    2916 23:16 00:01 4He target.
    2917 00:02 00:33
    2918 00:34 01:03
    2919 01:03 01:34
    2920 01:34 02:01
    2921 02:01 02:27 Run Stopped because of no beam.
    2922 02:35 03:02
    2923 03:03 03:30
    2924 03:31 04:00
    2925 04:00 04:32
    2926 04:54 05:21 40 uA on dummy target. BPMAY = 1.4 mm, 83k events. Probably junk.
    2927 05:23 05:37 40 uA on dummy target. Stopped to change prescales.
    2928 05:40 06:32 40 uA on dummy target.
    2929 junk
    2930 06:33 07:25 40 uA on dummy target.
    2931 07:38 07:45 4He target, junk, BPMA and BPMB y position changing.
    2932 07:45 ----- 4He target, 100 uA.